OpenStudy (anonymous):

Write an equation of the line with slope of -4/3 that passes thru the point (-6,2)?

OpenStudy (opcode):

Do you know about y=mx+b?

OpenStudy (opcode):

Where: M = Slope B = Y-intercept

OpenStudy (anonymous):

Yeah

OpenStudy (zale101):

Plug-in the values

OpenStudy (opcode):

Great just plug in the values. Slope: -4/3 x,y: (-6,2) So: \(y=mx+b\) \[2=\frac{-4}{3}(-6)+b\]

OpenStudy (opcode):

Can you solve for b or do you need my help? Do not be afraid to ask :-).

OpenStudy (anonymous):

Actually yeah if you wouldnt mind

OpenStudy (opcode):

Okay then. \[2=\frac{-4}{3}(-6)+b\] We need to solve for \(b\). So: \[2=\frac{-4}{3}(-6)+b\] Becomes: \[2=8+b\] Reverse the equality in \(2 = b+8\) in order to isolate b to the left hand side. \(2 = b+8\) is equivalent to \(b+8 = 2\): \(b+8 = 2\) Isolate terms with b to the left hand side. Subtract 8 from both sides: \(b+(8-8) = 2-8\) Look for two terms that sum to zero. \(8-8 = 0\) \(b = 2-8\) Evaluate \(2-8.\) \(2-8 = -6\) Answer: \(b = -6\)